OpenStudy (anonymous):

Find the exact value of the trigonometric function at the given real number. 1. cos (-13pi/20) 2. cos (10pi/3)

OpenStudy (anonymous):

for the first one i got \[-\sqrt{3}/2\]

OpenStudy (anonymous):

second one: 10pi/3 = 9pi/3 + pi/3 = 3pi + pi/3 = -cos(pi/3) = -1/2

OpenStudy (anonymous):

for the first one: -13pi/20 = -(10pi/20 + 3pi/20) = -pi/2 - 3pi/20 cos(-pi/2 - 3pi/20) = -sin(3pi/20) The exact value is -sin(3pi/20)

OpenStudy (anonymous):

well.. I just found out another exact value for the first equation: \[(-1/2)\sqrt{(1/2) (4-\sqrt{2(5-\sqrt{5}}))}\]

OpenStudy (anonymous):

I type in the first one wrong

OpenStudy (anonymous):

cos (-13pi/6)

OpenStudy (anonymous):

but following your steps should help me find the answer

OpenStudy (anonymous):

thats easy then. -13pi/6 = -12pi/6 -pi/6 = -2pi -pi/6 cos(-2pi -pi/6) = cos(-pi/6) = sqrt(3)/2

OpenStudy (awesome781):

Find the exact value of the trigonometric function at the given real number. Csc7pi/3

OpenStudy (awesome781):

Can someone help me on this question above